Find favg for the functions given on the interval and determine the value of c in the given

interval for which f(c) = favg.

a) f(x) = 9 − 2e^(4x+1) on [2,6]

b) 8 − cos (x/4) on [0 4π]


1
Expert's answer
2022-03-18T12:45:32-0400

Answer


We know that


"{f_{avg}} = \frac{1}{{b - a}}\int\limits_a^b {f\left( x \right)\,dx} \"


Solution (a) 


"f\left( x \right) = 9 - 2{e^{4x + 1}}" for the interval "[2, 6]"


"\begin{gathered}
 {f_{avg}} = \frac{1}{{6 - 2}}\int\limits_2^6 {\left( {9 - 2{e^{4x + 1}}} \right)dx} \\
 {f_{avg}} = \frac{9}{4}\int\limits_2^6 {\left( 1 \right)dx} - \frac{2}{4}\int\limits_2^6 {\left( {{e^{4x + 1}}} \right)dx} \\
 {f_{avg}} = \frac{9}{4}\left[ x \right]_2^6 - \frac{1}{2}\left[ {\frac{{{e^{4x + 1}}}}{4}} \right]_2^6 \\
 {f_{avg}} = \frac{9}{4}\left[ {6 - 2} \right] - \frac{1}{8}\left[ {{e^{4x + 1}}} \right]_2^6 \\
 {f_{avg}} = \frac{9}{4}\left[ 4 \right] - \frac{1}{8}\left[ {{e^{4\left( 6 \right) + 1}} - {e^{4\left( 2 \right) + 1}}} \right] \\
 {f_{avg}} = 9 - \frac{1}{8}\left[ {{e^{25}} - {e^9}} \right] \\
 {f_{avg}} = 9 - \frac{1}{8}\left[ {{e^{25}} - {e^9}} \right] \\
 {f_{avg}} = - 9000611395 \\ 
\end{gathered}"


Then for "f(c)=f_{avg}" means



"f\left( c \right) = {f_{avg}} \\
 9 - 2{e^{4c + 1}} = - 9000611395 \\
 2{e^{4c + 1}} = 9 + 9000611395 \\
 2{e^{4c + 1}} = 9000611404 \\
 {e^{4c + 1}} = 4500305702 \\
 4c + 1 = \ln \left( {4500305702} \right) \\
 4c + 1 = 22.22741117 \\
 4c = 21.22741117 \\
 c = 5.306852791 \\"



Solution (b) 


"f\left( x \right) = 8 - \cos \left( {\tfrac{x}{4}} \right)" for the interval "[0, 4\pi]"


"{f_{avg}} = \frac{1}{{4\pi - 0}}\int\limits_0^{4\pi } {\left( {8 - \cos \left( {\tfrac{x}{4}} \right)} \right)dx} \\
 {f_{avg}} = \frac{1}{{4\pi }}\int\limits_0^{4\pi } {\left( 8 \right)dx} + \frac{1}{{4\pi }}\int\limits_0^{4\pi } {\left( {\cos \left( {\tfrac{x}{4}} \right)} \right)dx} \\
 {f_{avg}} = \frac{8}{{4\pi }}\left[ x \right]_0^{4\pi } + \frac{1}{{4\pi }}\left[ {\frac{{\sin \left( {\tfrac{x}{4}} \right)}}{{\tfrac{1}{4}}}} \right]_0^{4\pi } \\
 {f_{avg}} = \frac{2}{\pi }\left[ {4\pi - 0} \right] + \frac{4}{{4\pi }}\left[ {\sin \left( {\tfrac{x}{4}} \right)} \right]_0^{4\pi } \\
 {f_{avg}} = 8 + \frac{1}{\pi }\left[ {\sin \left( {\tfrac{{4\pi }}{4}} \right) - \sin \left( {\tfrac{0}{4}} \right)} \right] \\
 {f_{avg}} = 8 + \frac{1}{\pi }\left[ {0 - 0} \right] \\
 {f_{avg}} = 8 \\"


Now for


"f\left( c \right) = {f_{avg}}\\
 8 - \cos \left( {\tfrac{c}{4}} \right) = 8 \\
 \cos \left( {\tfrac{c}{4}} \right) = 0 \\
 \frac{c}{4} = \frac{\pi }{2} \\
 c = 2\pi \\"